# A Self-Explainable Heterogeneous GNN for Relational Deep Learning

### v1

@@ 1 Introduction
Graph Neural Networks (GNNs) have increasingly become the de-facto standard for many predictive tasks involving networked data, such as physical systems [1, 2], Knowledge Graphs [3] and social networks [4]. By learning effective node embeddings, GNNs offer a unified framework for addressing various graph-based tasks, including node classification, graph classification, and link prediction. However, similar to other representation learning paradigms, GNNs often exhibit a black-box nature in their predictions. Numerous solutions have been proposed for post-hoc explainability of GNN predictions, primarily at the instance-based level [5, 6, 7, 8]. Yet, as with other deep learning architectures, the ability of these approaches to genuinely reflect the underlying reasoning of the predictor has been called into question [9]. To tackle this challenge, self-explainable GNNs [10, 11, 12] have recently emerged, aiming to ensure that GNN predictions are grounded in interpretable elements, such as subgraphs [13, 14] or prototypes [15, 16].
@@
Despite the widespread adoption of GNNs, most approaches are tailored for homogeneous graphs, where edge types are indistinguishable. While encoding edge types as features is a common workaround, the standard solution simply consists in concatenating the one-hot encoded edge type to node features, which eventually boils down to learning an edge-type specific bias. This limitation is particularly problematic in knowledge graphs, which typically feature numerous relations (corresponding to edge types), with only a few being pertinent to a specific predictive task—forming the so-called meta-paths. Existing approaches for heterogeneous GNNs either rely on domain experts to provide relevant meta-paths a priori [17, 18, 19, 20], or attempt to learn them from data by assigning different weights to various relations  [21, 22, 23, 24, 25, 26, 27, 28] , a solution that fails to scale with the number of candidate relations.
@@
Recently, MP-GNN [29] has been proposed as a solution for learning meta-paths without requiring user supervision. This approach employs a scoring function to predict the potential informativeness of partial meta-paths, enabling efficient exploration of the combinatorial space of candidate meta-paths. However, a key limitation of MP-GNN is its assumption that a node’s class is primarily dependent on the existence of a meta-path instance. While this assumption may be reasonable for knowledge graphs, it is unrealistic when dealing with relational databases, where entities are characterized by numerous categorical and numerical attributes. On the other hand, this same complexity makes relational databases particularly well-suited for GNN technology, as evidenced by the growing interest in what is now being termed relational deep learning [30].
@@
Figure 1: Left: Relational database schema for a medical domain. Right: Heterogeneous graph representation of (part of) the database. The highlighted subgraph shows a prototypical counts-of-counts pattern characterising positive patients, namely having at least two exempt prescriptions (indicated as true), each containing at least two medications. Existing heterogeneous GNNs struggle with these patterns as they need to learn a separate weight matrix for each edge type in the graph, while MPS-GNN is capable of learning them without any direct user supervision.
@@
In this paper, we extend the concept behind MP-GNN beyond the simple existential quantification of meta-paths. We introduce Meta-Path Statistics GNN (MPS-GNN), an approach that automatically identifies relevant meta-paths, where the informative content is determined by learnable statistics computed on their realizations. These include counts-of-counts statistics such as having at least two exempt prescriptions with at least two medications each to characterize patients with severe illness, as shown in Figure 1. An experimental evaluation on both synthetic and real-world relational database tasks demonstrate the significant advantages of the proposed solution over existing alternatives.
@@
Additionally, results show how the meta-path learning strategy behind MPS-GNN renders it inherently and genuinely self-explainable, in contrast to many existing self-explainable GNN architectures whose explanations often lack fidelity [11].
@@ 2 Related Work
Relational deep learning [30] has recently emerged as a paradigm advocating the application of deep learning technology, and GNNs in particular, to relational databases. The rationale behind this research direction is the popularity of relational databases as a mean to store relational information in a variety of application domains, combined with the fact that relational databases can be seen as heterogeneous graphs, with tables converted into sets of nodes and relations into (typed) edges between table entries [31]. This transformation allows the application of heterogeneous Graph Neural Networks (GNNs) to this kind of data.
@@
A common characteristics of most heterogeneous graphs, including those deriving from knowledge graphs and relational databases, is that only few relations convey relevant information when targeting a specific predictive task. For this reason, plain GNNs, that do not distinguish between edge types, struggle with these type of graphs. The most popular line of research for heterogeneous GNNs identifies meta-paths, i.e., sequences of relations, as primary sources of information. Existing approaches to incorporate meta-paths follow in two main categories. The former requires domain experts to identify relevant meta-paths for the task at hand [17, 18, 19, 20]. Clearly, this approach is suboptimal as it requires this domain information to be readily available. The alternative solutions, either utilize relation-specific graph convolutions, capturing relational patterns with distinct parameters or dedicated components for each type of relation  [21, 22, 24, 25] or focuses on graph transformation and multi-view learning to enhance relational representations [23, 26, 27, 28]. Although these methods are effective with a limited number of relations, their performance quickly deteriorates as the number of candidate relations grows.
@@
Recently, a novel approach named MP-GNN [29] has been introduced to tackle the aforementioned challenges and automatically learn relevant meta-paths from data. The approach leverages a scoring function predicting the potential informativeness of partial meta-paths to guide the search in the combinatorial space of candidate meta-paths. A major limitation of this approach is the fact that it assumes that the existential quantification of the meta-path is informative for the class label. This assumption makes the approach unsuitable for relational deep learning tasks, in which statistics extracted from table attributes are arguably crucial to characterize predictive targets. Our approach substantially generalizes the MP-GNN method, by designing a scoring function that can predict the informativeness of partial meta-paths in terms of the statistics that could be constructed on top of their realizations. This extension is crucial in allowing MPS-GNN to be effectively applied to relational deep learning settings, as shown by our experimental evaluation. More details about the comparison between MP-GNN and MPS-GNN in A.1.
@@
Explainability in GNN is a major research trend, with plenty of approaches for post-hoc explanation of GNN predictions (instance-based explainability [5, 6, 7, 8]) and others aiming to explain the GNN model as a whole (model-based explainability [32, 33, 34, 35]). Specialized metrics have been developed to estimate the faithfulness of an explanation in relation to the method’s input processing behavior [36, 37, 38, 39, 40]. Following a similar evolution in the XAI literature of convolutional neural networks, self-explainable GNNs [13, 14, 15, 16] have recently been proposed to encourage GNN models to adhere to their explanations by design. However, experimental studies have questioned the faithfulness of the explanations provided by these approaches [11, 41], highlighting the difficulty of achieving genuine explainability with GNNs. A key advantage of our proposed method is that meta-paths can naturally serve as model-level explanations, making MPS-GNN the first truly self-explainable GNN designed for relational deep learning applications. Our experimental evaluation confirms the faithfulness of the explanations to the model’s behaviour.
@@ 3 Preliminaries
A relational database $(\mathcal{T},\mathcal{L})$ consists of a collection of tables $\mathcal{T}=T_{1},...T_{n}$ and links between these tables $\mathcal{L}\subseteq T\times T$ Each table is a set $T=\{u_{1},...,u_{n}\}$ where the elements $u\in T$ are referred to as rows or entities. Each entity is a tuple $u=(\mathcal{P}_{u},\mathcal{K}_{u},a_{u})$ where $\mathcal{P}_{u}$ is the Primary Key that uniquely identifies the entity $u$; $\mathcal{K}_{u}$ is the set of Foreign Keys corresponding to a primary key in other tables, thus connecting the tables; $a_{u}$ corresponds to the Attributes of the entity $u$.
@@
From relational database to heterogeneous graph A relational database can be interpreted as an heterogeneous graph where row $u$ becomes node $v$; attributes $a_{u}$ become node features $x_{v}$; links $\mathcal{L}$ between entries of two tables are identified by the pair of primary $\mathcal{P}$ and foreign $\mathcal{K}$ keys in the two tables. Each pair of connected tables, originates a relation in the graph, specified by $r$.
@@ 4 Methodology
  • Greedy: a partially constructed meta-path $\xrightarrow{r_{1}}\xrightarrow{r_{2}}\ldots\xrightarrow{r_{i}}$ is extended by a next relation $\xrightarrow{r_{i+1}}$ without lookahead for possible completions $\xrightarrow{r_{i+1}}\ldots\xrightarrow{r_{L}}$. Similar as in [29], we try to estimate the potential informativeness of nodes reached by $\xrightarrow{r_{i+1}}$ by learned weights associated with the nodes. These weights represent putative features that can either be directly materialized as functions of the nodes’ attributes, or that can be constructed as features of meta-path extensions starting at the node.
@@
Note that both the greedy and local properties mirror core principles of growing decision trees, which are built incrementally, adding one relation at a time based on solving local classification sub-tasks.
@@ 4.1 Weighted multi-instance classification
$F({\mathcal{B}})=\sum_{v\in{\mathcal{B}}}\alpha(v,{\mathcal{B}})\cdot F(v),$ (1)
@@
$F(v,r,\Theta,{\boldsymbol{w}})=\Theta\cdot x_{v}\cdot\sum_{u\in\mathcal{N}^{r}_{v}}w_{u},$ (2)
where $x_{v}$ denotes the attribute vector of $v$, $\Theta$ is a trainable parameter vector, and ${\boldsymbol{w}}$ is a vector of trainable weights $w_{u}\in[0,1]$ assigned to $v$’s $r$-neighbors. The node feature function value for $v$ can, thus, be determined by $v$’s own attributes, or by aggregating the putative feature $w_{u}$ of its successors – or a combination of both elements.
@@
where $\sigma$ is the sigmoid function. In practice, we approximate (3) by a random sample of positive and negative bags.
@@ 4.2 Relation Scoring
The initial weighted multi-instance classification problem for our meta-path construction process is defined by letting each positive (negative) target node $v$ form a one-element bag $\{v\}$ with the corresponding label, and weight $\alpha(v,\{v\})=1$. Denote with ${\mathcal{S}}_{1}^{+},{\mathcal{S}}_{1}^{-}$ the sets of all initial positive and negative bags, respectively. At all iterations we select the relation that minimizes the loss
@@
If for some relation $r$ this loss is effectively minimized by suitable parameters $\Theta$, such that the $\sum w_{u}$ term in (2) plays no significant role, then the meta-path construction can terminate at this point (without adding $r$ to the meta-path). If, on the other hand, the $r$-successors and their weights ${\boldsymbol{w}}$ are needed to solve the current multi-instance classification task, then we need to extend the meta-path by $r$, and we need to capture the putative node features ${\boldsymbol{w}}$ by an actual node feature. A possible approach would be to set this up as a node regression task with target values $w_{u}$. However, due to the often very large set of alternative optimal solutions ${\boldsymbol{w}}$ in the minimization (4), this would lead to a too restrictive task. Our goal is to approximate the whole space of possible regression tasks defined by alternative ${\boldsymbol{w}}$ as a single weighted multi-instance classification task. For this, let $r_{i}$ denote the relation found to minimize (4) in iteration $i$ with optimal parameters $\Theta_{i}$. For each positive bag ${\mathcal{B}}^{+}_{i}\in\mathcal{S}_{i}^{+}$ define the new bag
${\mathcal{B}}^{+}_{i+1}=\cup_{v\in{\mathcal{B}}^{+}_{i}}\mathcal{N}^{r_{i}}_{v}$ (5)
@@
containing the $r_{i}$-children of the nodes in ${\mathcal{B}}^{+}_{i}$. Similarly for negative bags. For $u\in{\mathcal{B}}^{+}_{i+1}$ define the node weight
$\alpha(u,{\mathcal{B}}^{+}_{i+1})=\sum_{v\,:\,u\in\mathcal{N}^{r_{i}}_{v}}\Theta_{i}\cdot x_{v}\cdot\alpha(v,{\mathcal{B}}^{+}_{i}).$ (6)
Thus, nodes have a bigger weight in the subsequent multi-instance classification task, if they are $r_{i}$- successors of many nodes from the previous task, that make a significant contribution to the discriminant function (1) learned for the previous task. The sets of all ${\mathcal{B}}^{+}_{i+1}({\mathcal{B}}^{-}_{i+1})$ form the new training sets ${\mathcal{S}}^{+}_{i+1}({\mathcal{S}}^{-}_{i+1})$.

### v2

@@ 1 Introduction
Graph Neural Networks (GNNs) have increasingly become the de-facto standard for many predictive tasks involving networked data, such as physical systems (Sanchez-Gonzalez et al., 2018; Battaglia et al., 2016), Knowledge Graphs (Hamaguchi et al., 2017) and social networks (Wu et al., 2020). By learning effective node embeddings, GNNs offer a unified framework for addressing various graph-based tasks, including node classification, graph classification, and link prediction. However, similar to other representation learning paradigms, GNNs often exhibit a black-box nature in their predictions. Numerous solutions have been proposed for post-hoc explainability of GNN predictions, primarily at the instance-based level (Ying et al., 2019; Vu & Thai, 2020; Miao et al., 2022; Yuan et al., 2021). Yet, as with other deep learning architectures, the ability of these approaches to genuinely reflect the underlying reasoning of the predictor has been called into question (Longa et al., 2024). To tackle this challenge, self-explainable GNNs (Kakkad et al., 2023; Christiansen et al., 2023; Seo et al., 2023) have recently emerged, aiming to ensure that GNN predictions are grounded in interpretable elements, such as subgraphs (Wu et al., 2022; Yu et al., 2020) or prototypes (Zhang et al., 2022; Ragno et al., 2022).
@@
Despite the widespread adoption of GNNs, most approaches are tailored for homogeneous graphs, where edge types are indistinguishable. While encoding edge types as features is a common workaround, the standard solution simply consists in concatenating the one-hot encoded edge type to node features, which eventually boils down to learning an edge-type specific bias. This limitation is particularly problematic in knowledge graphs, which typically feature numerous relations (corresponding to edge types), with only a few being pertinent to a specific predictive task—forming the so-called meta-paths. Existing approaches for heterogeneous GNNs either rely on domain experts to provide relevant meta-paths a priori (Chang et al., 2022; Fu et al., 2020; Li et al., 2021; Wang et al., 2019), or attempt to learn them from data by assigning different weights to various relations  (Hu et al., 2020; Lv et al., 2021b; Mitra et al., 2022; Schlichtkrull et al., 2018; Yu et al., 2022; Yun et al., 2019b; 2022a; Zhu et al., 2019) , a solution that fails to scale with the number of candidate relations.
@@
Recently, MP-GNN (Ferrini et al., 2024) has been proposed as a solution for learning meta-paths without requiring user supervision. This approach employs a scoring function to predict the potential informativeness of partial meta-paths, enabling efficient exploration of the combinatorial space of candidate meta-paths. However, a key limitation of MP-GNN is the existential quantification assumption which states that a node’s class is primarily dependent on the existence of a meta-path instance, meaning just one instance is sufficient for accurate prediction.
its assumption that a node’s class is primarily dependent on the existence of a meta-path instance, meaning just one instance is sufficient for accurate prediction. While this assumption may be reasonable for knowledge graphs, it is unrealistic when dealing with relational databases, where entities are characterized by numerous categorical and numerical attributes. On the other hand, this same complexity makes relational databases particularly well-suited for GNN technology, as evidenced by the growing interest in what is now being termed relational deep learning (Fey et al., 2023).
@@
Figure 1: Left: Relational database schema for a medical domain. Right: Heterogeneous graph representation of (part of) the database. The highlighted subgraph shows a prototypical counts-of-counts pattern characterising positive patients, namely having at least two exempt prescriptions (represented by node feature T), each containing at least two medications. Existing heterogeneous GNNs struggle with these patterns as they need to learn a separate weight matrix for each edge type in the graph, while MPS-GNN is capable of learning the relevant meta path without any direct user supervision.
@@
In this paper, we extend the concept behind MP-GNN beyond the simple existential quantification of meta-paths. We introduce Meta-Path Statistics GNN (MPS-GNN), an approach that automatically identifies relevant meta-paths, where the informative content is determined by learnable statistics computed on their realizations. These include counts-of-counts statistics such as having at least two exempt prescriptions with at least two medications each to characterize patients with severe illness, as shown in Figure 1. Note that MP-GNN would fail to discriminate between the two patient nodes in the figure, as they both have at least one occurrence of the correct meta-path. An experimental evaluation on both synthetic and real-world relational database tasks demonstrate the significant advantages of the proposed solution over existing alternatives.
@@
Additionally, results show how the meta-path learning strategy behind MPS-GNN renders it inherently and genuinely self-explainable, in contrast to many existing self-explainable GNN architectures whose explanations often lack fidelity (Christiansen et al., 2023).
@@ 2 Related Work
Relational deep learning (Fey et al., 2023) has recently emerged as a paradigm advocating the application of deep learning technology, and GNNs in particular, to relational databases. The rationale behind this research direction is the popularity of relational databases as a mean to store relational information in a variety of application domains, combined with the fact that relational databases can be seen as heterogeneous graphs, with tables converted into sets of nodes and relations into (typed) edges between table entries (Robinson et al., 2024). This transformation allows the application of heterogeneous Graph Neural Networks (GNNs) to this kind of data.
@@
A common characteristics of most heterogeneous graphs, including those deriving from knowledge graphs and relational databases, is that only few relations convey relevant information when targeting a specific predictive task. For this reason, plain GNNs, that do not distinguish between edge types, struggle with these type of graphs. The most popular line of research for heterogeneous GNNs identifies meta-paths, i.e., sequences of relations, as primary sources of information. Existing approaches to incorporate meta-paths follow in two main categories. The former requires domain experts to identify relevant meta-paths for the task at hand (Chang et al., 2022; Fu et al., 2020; Li et al., 2021; Wang et al., 2019). Clearly, this approach is suboptimal as it requires this domain information to be readily available. The alternative solutions, either utilize relation-specific graph convolutions, capturing relational patterns with distinct parameters or dedicated components for each type of relation  (Hu et al., 2020; Lv et al., 2021b; Schlichtkrull et al., 2018; Yu et al., 2022) or focuses on graph transformation and multi-view learning to enhance relational representations (Mitra et al., 2022; Yun et al., 2019b; 2022a; Zhu et al., 2019). Although these methods are effective with a limited number of relations, their performance quickly deteriorates as the number of candidate relations grows.
@@
Recently, a novel approach named MP-GNN (Ferrini et al., 2024) has been introduced to tackle the aforementioned challenges and automatically learn relevant meta-paths from data. The approach leverages a scoring function predicting the potential informativeness of partial meta-paths to guide the search in the combinatorial space of candidate meta-paths. A major limitation of this approach is the fact that it assumes that the existential quantification of the meta-path is informative for the class label. This assumption makes the approach unsuitable for relational deep learning tasks, in which statistics extracted from table attributes are arguably crucial to characterize predictive targets. Our approach substantially generalizes the MP-GNN method, by designing a scoring function that can predict the informativeness of partial meta-paths in terms of the statistics that could be constructed on top of their realizations. This extension is crucial in allowing MPS-GNN to be effectively applied to relational deep learning settings, as shown by our experimental evaluation. More details about the comparison between MP-GNN and MPS-GNN in 4.6.
@@
Explainability in GNN is a major research trend, with plenty of approaches for post-hoc explanation of GNN predictions (instance-based explainability (Ying et al., 2019; Vu & Thai, 2020; Miao et al., 2022; Yuan et al., 2021)) and others aiming to explain the GNN model as a whole (model-based explainability (Chen et al., 2024; Wang & Shen, 2022; Azzolin et al., 2022; Yuan et al., 2020)). Specialized metrics have been developed to estimate the faithfulness of an explanation in relation to the method’s input processing behavior (Agarwal et al., 2023; Yuan et al., 2022; Amara et al., 2022; 2023; Zheng et al., 2023). Following a similar evolution in the XAI literature of convolutional neural networks, self-explainable GNNs (Wu et al., 2022; Yu et al., 2020; Zhang et al., 2022; Ragno et al., 2022) have recently been proposed to encourage GNN models to adhere to their explanations by design. However, experimental studies have questioned the faithfulness of the explanations provided by these approaches (Christiansen et al., 2023; Azzolin et al., 2025), highlighting the difficulty of achieving genuine explainability with GNNs. A key advantage of our proposed method is that meta-paths can naturally serve as model-level explanations, making MPS-GNN the first truly self-explainable GNN designed for relational deep learning applications. Our experimental evaluation confirms the faithfulness of the explanations to the model’s behaviour.
@@ 3 Preliminaries
A relational database $(\mathcal{T},\mathcal{L})$ consists of a collection of tables $\mathcal{T}=T_{1},...T_{n}$ and links between these tables $\mathcal{L}\subseteq T\times T$ Each table is a set $T=\{e_{1},...,e_{n}\}$ where the elements $e\in T$ are referred to as rows or entities. Each entity is a tuple $e=(\mathcal{P}_{e},\mathcal{K}_{e},a_{e})$ where $\mathcal{P}_{e}$ is the Primary Key that uniquely identifies the entity $e$; $\mathcal{K}_{e}$ is the set of Foreign Keys corresponding to a primary key in other tables, thus connecting the tables; $a_{e}$ corresponds to the Attributes of the entity $e$.
@@
From relational database to heterogeneous graph A relational database can be interpreted as an heterogeneous graph where row $e$ becomes node $v$; attributes $a_{e}$ become node features $x_{v}$; links $\mathcal{L}$ between entries of two tables are identified by the pair of primary $\mathcal{P}$ and foreign $\mathcal{K}$ keys in the two tables. Each pair of connected tables, originates a relation in the graph, specified by $r$.
@@ 4 Methodology
Figure 2: Outline of greedy local meta-path construction
@@
  • Greedy: a partially constructed meta-path $\xrightarrow{r_{1}}\xrightarrow{r_{2}}\ldots\xrightarrow{r_{i}}$ is extended by a next relation $\xrightarrow{r_{i+1}}$ without lookahead for possible completions $\xrightarrow{r_{i+2}}\ldots\xrightarrow{r_{L}}$. Similar as in Ferrini et al. (2024), we try to estimate the potential informativeness of nodes reached by $\xrightarrow{r_{i+1}}$ by learned weights associated with the nodes. These weights represent putative features that can either be directly materialized as functions of the nodes’ attributes, or that can be constructed as features of meta-path extensions starting at the node.
@@
Figure2 illustrates the high-level principles of our approach. Given an initial (binary) node classification task, a first relation is identified that could solve the task with the help of a putative node feature (weight) $\boldsymbol{w}$ on the successor nodes. Then a surrogate classification task is set up whose target is to materialize the putative feature as a feature computable from the data. This surrogate task takes the form of a weighted multi-instance classification task, which can be seen as an abstraction of a direct regression problem with target $\boldsymbol{w}$ (see Section 4.2).
Note that both the greedy and local properties mirror core principles of growing decision trees, which are built incrementally, adding one relation at a time based on solving local classification sub-tasks. The improvements in time complexity with respect to an algorithm that looks at every possible relations in the meta-path construction are detailed in section 4.2.1. In Section 4.1, we describe the weighted multi-instance classification task and its application in our scenario. Section 4.2 outlines the methodology for constructing meta-paths by scoring graph relations, including the examples based on Figure 1. In Section 4.3, we detail the GNN used in our setup, and Section 4.4 presents the complete framework of the proposed approach.
@@ 4.1 Weighted multi-instance classification
$F({\mathcal{B}})=\sum_{v\in{\mathcal{B}}}\alpha(v,{\mathcal{B}})f(v),$ (1)
@@
$f(v,r,\Theta,{\boldsymbol{w}})=\left\{\begin{array}[]{ll}\Theta^{T}x_{v}&\mbox{if}\ \mathcal{N}^{r}_{v}=\emptyset\\ \Theta^{T}x_{v}\sum_{u\in\mathcal{N}^{r}_{v}}w_{u}&\mbox{if}\ \mathcal{N}^{r}_{v}\neq\emptyset\end{array}\right.$ (2)
where $x_{v}$ denotes the attribute vector of $v$, $\Theta$ is a trainable parameter vector, and ${\boldsymbol{w}}$ is a vector of trainable weights $w_{u}\in[0,1]$ assigned to $v$’s $r$-neighbors. The node feature function is thus computed as a combination of $v$’s own attributes and the putative feature $w_{u}$ of its successors. When the node attributes $x_{v}$ are not informative for solving the multi-instance classification task, then parameters $\Theta$ can be learned that make $\Theta^{T}x_{v}$ constant for all $v$, and thus this part of the node feature function becomes irrelevant (a suitable $\Theta$ exists e.g. under the mild assumption that $x_{v}$ contains at least one categorical attribute  [^1] In a heterogeneous graph built from a relational database, this would be the table the node belongs to. If node representations are simply embeddings, one can achieve the same goal by concatenating a constant feature to the embedding. in a one-hot encoding: then $\Theta^{T}x_{v}=1$ for $\Theta$ that is set to 1 for all entries corresponding to the one-hot encoding of the categorical attribute, and zero everywhere else). The factor $\sum_{u\in\mathcal{N}^{r}_{v}}w_{u}$ captures a dependence of $f(v)$ on the $r$-neighborhood of $v$. When the $r$-neighborhood is non-informative because neither the number nor the identity of nodes $u\in\mathcal{N}^{r}_{v}$ has discriminative value, then the $\sum_{u\in\mathcal{N}^{r}_{v}}w_{u}$ factor can be made irrelevant by learning constant weights $w_{u}$.
@@
where $\sigma$ is the sigmoid function. In practice, given that the number of terms in the sum is quadratic in the number of training examples, we approximate (3) by a random sample of positive and negative bags.
@@ 4.2 Relation Scoring
The initial weighted multi-instance classification problem for our meta-path construction process is defined by letting each positive (negative) target node $v$ form a one-element bag $\{v\}$ with the corresponding label, and weight $\alpha(v,\{v\})=1$. Denote with ${\mathcal{S}}_{1}^{+},{\mathcal{S}}_{1}^{-}$ the sets of all initial positive and negative bags, respectively. At all iterations we select the relation that minimizes the loss (also referred to as the scoring function)
@@
If all candidate relations fail to minimize the loss, i.e., optimizing $\Theta,{\boldsymbol{w}}$ does not lead to substantially lower loss than using random parameters (i.e., does not improve by at least 30%), then the meta-path construction terminates and the current meta-path is returned. Otherwise, the current meta-path is extended with the minimal loss relation $r$. At this point, the problem becomes capturing the putative node features ${\boldsymbol{w}}$ by actual features.
@@
A possible approach would be to set this up as a node regression task with target values $w_{u}$. However, due to the often very large set of alternative optimal solutions ${\boldsymbol{w}}$ in the minimization (4), this would lead to a too restrictive task. Our goal is to approximate the whole space of possible regression tasks defined by alternative ${\boldsymbol{w}}$ as a single weighted multi-instance classification task. For this, let $r_{i}$ denote the relation found to minimize (4) in iteration $i$ with optimal parameters $\Theta_{i}$. For each positive bag ${\mathcal{B}}^{+}\in\mathcal{S}_{i}^{+}$ define the new bag
${\mathcal{B}}^{+}_{\emph{new}}=\cup_{v\in{\mathcal{B}}^{+}}\mathcal{N}^{r_{i}}_{v}$ (5)
@@
containing the $r_{i}$-children of the nodes in ${\mathcal{B}}^{+}$. Similarly for negative bags. For $u\in{\mathcal{B}}^{+}_{\emph{new}}$ define the node weight by the following sum over all nodes $v\in\mathcal{B}^{+}$ that have $u$ as an $r_{i}$-neighbor:
$\alpha(u,{\mathcal{B}}^{+}_{\emph{new}})=\sum_{\{v\in\mathcal{B}^{+}\,:\,u\in\mathcal{N}^{r_{i}}_{v}\}}\Theta_{i}^{T}x_{v}\alpha(v,{\mathcal{B}}^{+}).$ (6)
This definition of the weights enables to essentially ignore in the setup of the multi-instance classification task for iteration $i+1$ those nodes $u$ that are $r_{i}$-successors only of nodes $v$ that did not play a major role in solving the task of the previous iteration – either because of a low absolute value of $\Theta_{i}^{T}x_{v}$, or because $\alpha(v,{\mathcal{B}}^{+}_{i})$ already had a low absolute value. The sets of all ${\mathcal{B}}^{+}_{\emph{new}}({\mathcal{B}}^{-}_{\emph{new}})$ form the new training sets ${\mathcal{S}}^{+}_{i+1}({\mathcal{S}}^{-}_{i+1})$.
Toy Example:
To provide a clear understanding of the proposed approach, we illustrate it with the toy example of Figure 1, where patients are positive if, and only if, they have at least two exempt prescriptions, each containing at least two medications.
Initially, the approach evaluates the potential informativeness of two relations originating from patients: relation $a$ and relation $b$. Figure 3(a) illustrates the scoring process for relation a. The first step involves instantiating the node feature functions for patient nodes 0 and 1 by applying Equation 2. Since both nodes share the same unique neighbor, the loss function 4 reduces to the constant $\frac{1}{2}$, which cannot be minimized by $\Theta$ and ${\boldsymbol{w}}$.
(a) Scoring relation $a$
(b) Scoring relation $b$
Figure 3: Scoring the first two relations
Similarly, relation $b$ is scored. The loss function 4 now becomes
$\sigma(\Theta^{T}x_{1}(w_{4}+w_{5}+w_{6})-\Theta^{T}x_{0}(w_{2}+w_{3}+w_{4})),$
where $x_{0}=x_{1}=(1,0)^{T}$ is the attribute vector representing the T value in a one-hot encoding. This loss can be brought arbitrarily close to zero, e.g. by $w_{6}\rightarrow 0$, $w_{5}\rightarrow 0$, $w_{3}\rightarrow 1$, $w_{2}\rightarrow 1$, and $\Theta=(Z,0)^{T}$ with $Z\gg 0$. Consequently, relation $b$ scores higher than relation $a$ and is selected for further extension. Note that the low loss of relation $b$ is entirely due to its potential informativeness: the $b$ neighborhoods of the nodes 0 and 1 are completely isomorphic, and therefore the meta path consisting of $b$ alone provides no discriminative information.
To extend the meta path prefix $b$ we set up the weighted multi-instance classification task for the second iteration. This gives us a positive bag $\mathcal{B}_{2}$ containing nodes $\{2,3,4\}$ and a negative bag $\mathcal{B}_{3}$ containing nodes $\{4,5,6\}$. Node weights for the bags are computed according to Equation (6), and here simply yield uniform weights $\alpha(k,\mathcal{B}_{2})=Z$ for all $k\in\{2,3,4\}$, and $\alpha(k,\mathcal{B}_{3})=Z$ for all $k\in\{4,5,6\}$. Note that node $4$ has separate weights for the two bags it is part of.
Figure 4: Bag generation and scoring of relation $c$.
For solving the new classification task we score the candidate relations $c$ and $d$. The right part of Figure 4 illustrates the scoring of $c$. Due to the indistinguishable structure of the $c$-neighborhood for bags $\mathcal{B}_{2}$ and $\mathcal{B}_{3}$ the loss function 4 now reduces again to the constant $1/2$, as in the scoring of relation $a$.
H Refer to caption
Figure 5: Scoring relation $d$.
For scoring the relation $d$ we obtain as the loss function 4
$\sigma\left(Z\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}(w_{12}-w_{10}-w_{11})+Z\Theta^{T}\begin{bmatrix}0\\ 1\end{bmatrix}(w_{10}+w_{11}-1)\right)$